### v1

# Mamba-FCS: Joint Spatio-Frequency Feature Fusion, Change-Guided Attention, and SeK Loss for Enhanced Semantic Change Detection in Remote Sensing

Abstract
Semantic Change Detection (SCD) from remote sensing imagery requires models balancing extensive spatial context, computational efficiency, and sensitivity to class-imbalanced land-cover transitions. While Convolutional Neural Networks excel at local feature extraction but lack global context, Transformers provide global modeling at high computational costs. Recent Mamba architectures based on state-space models offer compelling solutions through linear complexity and efficient long-range modeling. In this study, we introduce Mamba-FCS, a SCD framework built upon Visual State Space Model backbone incorporating, a Joint Spatio-Frequency Fusion block incorporating log-amplitude frequency domain features to enhance edge clarity and suppress illumination artifacts, a Change-Guided Attention (CGA) module that explicitly links the naturally intertwined BCD and SCD tasks, and a Separated Kappa (SeK) loss tailored for class-imbalanced performance optimization. Extensive evaluation on SECOND and Landsat-SCD datasets shows that Mamba-FCS achieves state-of-the-art metrics, 88.62% Overall Accuracy, 65.78% $F_{\mathrm{scd}}$, and 25.50% SeK on SECOND, 96.25% Overall Accuracy, 89.27% $F_{\mathrm{scd}}$, and 60.26% SeK on Landsat-SCD. Ablation analyses confirm distinct contributions of each novel component, with qualitative assessments highlighting significant improvements in SCD. Our results underline the substantial potential of Mamba architectures, enhanced by proposed techniques, setting a new benchmark for effective and scalable semantic change detection in remote sensing applications. The complete source code, configuration files, and pre-trained models will be publicly available upon publication.
Index Terms:
Semantic change detection, Remote sensing imagery, State-space models, Spatial–frequency fusion, Separated Kappa
I Introduction
Change Detection (CD) is a widely studied and increasingly popular field in remote sensing, which involves identifying alterations in the Earth’s surface by comparing remote sensing images of the same area captured at different times[1]. CD plays a vital role in various applications, including urban planning, land cover and land use (LCLU) analysis, disaster assessment, ecosystem monitoring, and natural resource management[2, 3, 4, 5].
CD tasks are primarily categorized into Binary Change Detection (BCD) and Semantic Change Detection (SCD), which are inherently coupled[6, 7]. BCD focuses on detecting whether a change has occurred or not between remote sensing images, classifying areas as either ”change” or ”no change.” In contrast, SCD goes a step further by identifying the specific nature of the change, providing detailed “from-to” transition information between different land cover or land use types[8].
Furthermore, CD Algorithms can be further divided into supervised and unsupervised approaches. Unsupervised methods such as K-means, ISODATA and graph cut methods aim to identify change labels without the reliance on training with labeled data[9, 10, 11]. However, this tends to limit the performance especially in complex semantic transitions. Supervised methods, on the contrary, leverage labeled data to learn richer representations of these complex transitions in order to gain superior performance[12].
Early CD algorithms relied on image differencing, change-vector analysis and Bayesian fusion of pre/post semantic maps[13]. These multi-stage heuristics accumulate errors and struggle with spectral inconsistencies caused by varying illumination, phenology or sensor viewing angles[14].
Traditional methods attempted to mitigate these spectral inconsistencies using physical or empirical correction methods. Methods such as C-Correction and Minnaert correction as well as Normalization methods are examples of this[15]. While these approaches can improve robustness of traditional methods induced due to variability, they usually rely on additional data such as elevation models or require parameter tuning[16]. These shortcomings highlight the need for data-driven alternatives for CD.
Introduction of deep learning architectures opened a new avenue for addressing most of these challenges. Early end‐to‐end designs based on convolutional neural networks (CNNs) overcame many pitfalls of handcrafted pipelines, yet their limited receptive field still hampers the capture of long‐range context in large, heterogeneous scenes[17, 18, 19].
The Attention mechanism, in Transformer Models[20] for incorporating long-range dependencies in sequence modeling, was adopted by Vision Transformers(ViT) for modeling global dependencies in computer vision tasks. These integrate windowed attentions as well as global attention mechanisms for better modeling of global dependencies. These methods and variants have successfully been used in CD Algorithms with competitive results[21]. However, their quadratic complexity renders dense, high‐resolution prediction prohibitively expensive.
State‐space models (SSMs) such as S4[22] and, more recently, Mamba[23] offer an attractive middle ground- they deliver global sequence modelling with linear complexity and hardware‐friendly implementations. These models have since been adopted for computer vision tasks [24] and later for CD [25, 26] with promising performance.
Despite these advances, real‑world deployment still suffers from class imbalance, isolation of semantic and change detection branches, and neglect of frequency domain cues. In particular, because certain semantic transitions often occupy a small fraction of pixels in certain datasets, conventional models tend to be biased toward more prevalent classes[27]. Furthermore, although SCD and BCD are naturally interconnected, existing approaches like [28, 29] keep change and semantic decoders isolated, preventing mutual reinforcement and resulting in blurred boundaries and inconsistent bi‐temporal predictions. Moreover, although Fourier‐domain representations are proven to suppress illumination artifacts and emphasize genuine structural differences in BCD[30], existing Mamba frameworks for SCD predominantly focus on spatial features, with limited exploration of frequency-domain fusion benefits.
Furthermore, while the Separated Kappa (SeK) coefficient has become the standard metric for evaluating SCD performance since its introduction on the SECOND benchmark [31], its use as a minimizable loss function remains unexplored.
In this work, we introduce Mamba-FCS, a novel approach that integrates frequency-domain feature fusion and change-guided attention within a Mamba backbone to address limitations of purely spatial processing, and decoder isolation. By incorporating SeK as a loss term, we emphasis semantic consistency within detected change regions. Extensive experiments on the SECOND and LandSat-SCD benchmarks demonstrate substantial improvements in detecting non-salient changes, enhancing recall for rare transitions, and refining semantic boundaries. In summary, our main contributions includes:
  • Proposing Joint Spatio-Frequency Fusion Mechanism, that effectively combines spatial context with frequency-domain information across multiple stages, improving feature representation,
  • Proposing a Change-Guided Attention (CGA) module that iteratively incorporates an intermediate change-probability map into each semantic decoder stage, enabling mutual reinforcement between BCD and SCD branches for higher-fidelity change masks and sharper semantic boundaries, and
  • Proposing the SeK Loss, leveraging the SeK metric as a loss function to directly optimize for semantic consistency within detected change regions, mitigating the impact of class imbalance.
The remainder of this paper is organized as follows. Section II reviews related work and motivates our contributions. Section III presents the Mamba-FCS framework, detailing the key components including Joint Spatio-Frequency Fusion block, Change-Guided Attention module, and Separated Kappa loss formulation. Section IV describes the experimental methodology and datasets. Section V presents comprehensive results and ablation analyses. Finally, Section VI draws conclusions.
II Related Work
In this section, we review deep learning-based CD methods for high-resolution optical remote sensing imagery, organized into CNN-based, Transformer-based, State Space Model (SSM)-based, and Frequency-Domain-based approaches.
II-A CNN-Based Methods
CNNs have been pivotal in early deep learning approaches for CD due to their robust ability to extract local features. Daudt et al. in [32] proposed the first fully convolutional networks (FCNs) for BCD task, with three variants: FC-Siam-Conc, FC-EF, and FC-Siam-Diff. These models employed a Siamese CNN architecture to process bi-temporal inputs effectively. Subsequent advancements built upon this foundation. For instance, Fang et al. in [33] developed a Siamese-based architecture with a densely connected CNN to enhance feature interaction between bi-temporal images. Similarly, Zhang et al. in [34] introduced multi-level, fine-grained detection by applying deep supervision to differential features extracted at various CNN stages. Shi et al. in [35] further refined this approach by incorporating an attention mechanism to capture more discriminative features. Zhao et al. in [36] investigated the impact of different fusion strategies on BCD performance, proposing a dual encoder-decoder architecture to boost detection accuracy. More recently, Chen et al. in [37] devised an unsupervised sample generation method by swapping image patches, enabling change detection training on single-temporal images.
In contrast to BCD, SCD presents greater challenges but holds significant practical value [38, 39]. Early CNN-based SCD methods adopted a multitask learning framework to simultaneously predict binary change maps and semantic change categories [39]. Ding et al. in [29] proposed an advanced CNN architecture to address the limited information exchange between semantic and change decoders, thereby improving SCD performance.
Despite these advancements, CNN-based models exhibit limitations in CD tasks. CNN architectures suffer from restricted receptive fields, which hinder their ability to capture long-range contextual information essential for detecting semantic changes across large regions. This limitation has driven research toward Transformer-based methods, which offer a global view of the scene [40].
II-B Transformer-Based Methods
With the introduction of Vision Transformers (ViT)[20] in capturing long-range dependencies in images, CD research has increasingly adopted transformer architectures. For BCD, Chen et al. [41] were among the first to apply transformers by converting multi-temporal images into semantic tokens, thereby modeling spatial-temporal relationships to enhance detection accuracy. Bandara et al. in [21] further advanced this by proposing a pure transformer-based Siamese network, which employs a Siamese transformer encoder paired with a multilayer perceptron (MLP) decoder, eliminating the need for convolutional feature extractors.
In the context of SCD, Ding et al.[42] introduced the semantic change transformer, an adaptation of the CSWin Transformer, designed to explicitly learn semantic transitions. This model features a triple encoder-decoder architecture that enhances spatial features and integrates spatio-temporal dynamics with task-specific prior information to reduce learning disparities. Similarly, SMBCNet[43] leverages a transformer-based architecture to perform change detection through semantic segmentation. By integrating a cross-scale enhancement module and a multi-branch change fusion module, SMBCNet effectively captures global information and handles diverse change types. Additionally, STGNet [8] addresses SCD by guiding multitask learning through spatio-temporal semantic interaction, enhancing spatial details and employing a bidirectional guidance module to improve feature extraction in complex scenes.
Despite the advancements brought by transformer-based methods, their computational complexity, primarily due to the quadratic complexity of self-attention mechanisms, poses challenges for large-scale remote sensing datasets. To address this, recent research has explored alternative architectures such as SSMs[22].
Figure 1: Overview of the proposed Mamba-FCS architecture, featuring a VSSM-based encoder and decoder for three tasks- BCD map and SCD maps for $T_{1}$ and $T_{2}$. The encoder processes pre ($I^{T_{1}}$) and post ($I^{T_{2}}$) images through four hierarchical stages (Stage I to Stage IV), extracting multi-scale feature maps. These are fed into the binary change decoder and semantic map decoders, with binary change decoder’s stage-wise outputs guiding the semantic decoders.
II-C SSM-Based Methods
SSMs, particularly the Mamba architecture, have emerged as a promising alternative to Transformers due to their linear computational complexity and ability to model long-range dependencies [23]. In computer vision, the introduction of Visual State Space Model (VMamba) [44] and Vision Mamba [24] has facilitated their adoption for change detection tasks.
For BCD, [28] proposed MambaBCD, which leverages the VMamba architecture to extract comprehensive global spatial features from bi-temporal images. This model uses a change decoder with spatio-temporal relationship modeling, achieving superior performance. Similarly, CDMamba [26] introduces the Scaled Residual ConvMamba (SRCM) block, combining Mamba’s global feature extraction with convolutional layers for enhanced local detail capture. An Adaptive Global Local Guided Fusion (AGLGF) block facilitates bi-temporal feature interaction, improving detection accuracy. MSCNet [45]employs a Mamba-based self-correction network with a spatial-channel interaction fusion architecture, using a momentum update strategy to generate pseudo-labels and correct manual label inaccuracies, enhancing performance on challenging BCD tasks. Additionally, the Iterative Mamba Diffusion Change-Detection Model (IMDCD) [46] integrates Mamba with diffusion models to iteratively refine change detection maps, increasing sensitivity to subtle changes in complex scenes.
For SCD, MambaSCD [28] adapts the Visual Mamba framework to model complex multi-temporal relationships, effectively capturing semantic transitions.
However, above Mamba-based methods primarily operate in the spatial domain, often neglecting frequency-domain information that could suppress illumination artifacts and highlight structural differences in multi-temporal imagery.
II-D Frequency‐Domain–Based Methods
Frequency-domain methods have emerged as a robust approach for remote sensing CD by transforming images from the spatial domain to the frequency domain, typically using the Fourier transform. This transformation highlights structural differences, such as edges and textures, while suppressing illumination artifacts common in multi-temporal imagery [47]. These methods are particularly effective for detecting changes in texture or periodic patterns and for handling multimodal data from diverse sensors, such as optical and synthetic aperture radar (SAR).
For BCD, [48] proposed MFGFNet, which employs multiple global filters in the frequency domain to enhance boundary sharpness and preserve edge information in change regions. By integrating frequency-domain processing with multi-scale feature extraction, MFGFNet improves detection accuracy on datasets like LEVIR-CD. Similarly, the Frequency-Enhanced Mamba for Remote Sensing Change Detection (FEMCD) [30] introduces a difference-guided state-space model (DGSSM) to extract change-related features and a DCT-aided Mamba decoder (DCTMD) to refine minor and texture changes using frequency information, achieving superior BCD performance. SpectMamba [49] further enhances BCD by integrating frequency information with the Mamba architecture to address high-frequency subtle changes and periodic structural changes in multispectral remote sensing images, demonstrating improved performance.
Despite these advancements, to the best of our knowledge, existing literature has not yet extensively explored integrating frequency and spatial information with the Mamba architecture for SCD. Additionally, current Mamba-based methods for SCD often employ isolated decoders, potentially limiting the mutual reinforcement that could benefit both tasks. Moreover, these methods have yet to explore the use of SeK metric as minimizable loss.
Figure 2: Architecture of the Visual State-Space Model (VMamba) backbone, used as the shared encoder. Following a patch-partition layer, four stages of Visual State-Space (VSS) blocks employ 2D selective scanning to progressively downsample spatial resolution (from $H,W$ to $H/32,W/32$) while increasing channel width. Feature shapes extracted from each stage are annotated adjacent to the blocks.
III Methodology
In this section, we first present an overview of the proposed architecture. We then describe each component of the architecture in detail. Finally, we discuss the loss functions employed in our methodology.
III-A Overview of the Network Architecture
The proposed Mamba-FCS framework utilizes a Siamese network architecture with a shared encoder for processing bi-temporal imagery, followed by three decoders to jointly perform SCD and BCD. As illustrated in Figure 1, the framework processes input image pairs $I^{T_{1}},I^{T_{2}}\in\mathbb{R}^{3\times H\times W}$, where 3 denotes the RGB color channels, and $H$ and $W$ represent the image height and width respectively, corresponding to pre-change and post-change temporal scenes. These bi-temporal inputs are processed by a shared Visual State Space Model (VSSM)-based backbone encoder, denoted as $\mathcal{F}_{encoder}$, which is detailed in Section III-B. This encoder, as seen in figure 2 extracts multi-scale feature maps at four hierarchical levels, providing a robust foundation for subsequent decoding and fusion operations. The feature extraction process is formalized as follows.
$X_{1}^{T_{1}},X_{2}^{T_{1}},X_{3}^{T_{1}},X_{4}^{T_{1}}=\mathcal{F}_{encoder}(I^{T_{1}})$ (1)
$X_{1}^{T_{2}},X_{2}^{T_{2}},X_{3}^{T_{2}},X_{4}^{T_{2}}=\mathcal{F}_{encoder}(I^{T_{2}})$
where $X_{i}^{T_{1}},X_{i}^{T_{2}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ denote the feature maps extracted from the $i$-th stage of the encoder for time steps $T_{1}$ and $T_{2}$, respectively, with $C_{i}$, $H_{i}$, and $W_{i}$ representing the channel depth, height, and width at each stage.
These multi-scale feature maps are then fed to the three decoder networks as seen in figure 1. First, our Binary Change Decoder, $\mathcal{F}_{BCD}$, detailed in section III-D, detects changes between the two time steps by fusing corresponding feature maps from $I^{T_{1}}$ and $I^{T_{2}}$ at each stage. The fusion is performed using Joint Spatio-Frequency Feature Fusion Mechanism $\mathcal{F}_{fusion}$ as described in section III-C. As illustrated in figure 4, the Binary Change Decoder employs a hierarchical refinement strategy to generate the final binary change detection output $Y_{BCD}\in\mathbb{R}^{2\times H\times W}$, where the two channels correspond to change and no-change classes. Additionally, the decoder produces intermediate change maps $CM_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ for $i\in\{1,2,3,4\}$ representing each hierarchical level. This can be formulated as,
$Y_{\text{BCD}},\{CM_{i}\}_{i=1}^{4}=\mathcal{F}_{BCD}(\{X_{i}^{T_{1}},X_{i}^{T_{2}}\}_{i=1}^{4})$ (2)
The change maps $\{CM_{i}\}_{i=1}^{4}$ will be used in CGA Module in Semantic Change Detection, detailed in the section III-E1.
For SCD, the framework incorporates two dedicated Semantic Map Decoders, $\mathcal{F}_{SCD}^{T_{1}}$ and $\mathcal{F}_{SCD}^{T_{2}}$, one for each time step as detailed in section III-E. These decoders, as illustrated in figure 5, utilize encoder extracted features $\{X_{i}^{T_{j}}\}_{i=1}^{4}$ ($j\in\{1,2\}$) and intermediate change maps $\{CM_{i}\}_{i=1}^{4}$ to enhance semantic segmentation by prioritizing changed regions. The process can be defined as,
$Y^{T_{1}}=\mathcal{F}_{SCD}^{T_{1}}(\{X_{i}^{T_{1}}\}_{i=1}^{4},\{CM_{i}\}_{i=1}^{4})$ (3)
$Y^{T_{2}}=\mathcal{F}_{SCD}^{T_{2}}(\{X_{i}^{T_{2}}\}_{i=1}^{4},\{CM_{i}\}_{i=1}^{4})$
where $Y^{T_{1}},Y^{T_{2}}\in\mathbb{R}^{N\times H\times W}$ are the semantic segmentation maps at $T_{1}$ and $T_{2}$, respectively, with $N$ channels corresponding to $N$ semantic classes.
Figure 3: The Joint Spatio-Frequency Feature Fusion ($F_{\text{fusion}}$) block, integrated at each decoder stage (left panel), concatenates spatial features $X^{T_{1}}_{i}$, $X^{T_{2}}_{i}$, log-amplitude frequency-domain features $F^{T_{1}}_{i}$, $F^{T_{2}}_{i}$, and the absolute difference map $D_{i}$. These are compressed and refined by a channel-then-spatial CBAM module (right panel) to produce the fused tensor $X^{\text{fused}}_{i}$.
III-B VSSM Backbone Encoder- VMamba
The backbone encoder in the proposed Mamba-FCS framework is based Visual State Space Models (VMamba) [44], a state-of-the-art vision backbone designed for efficient visual representation learning with linear computational complexity. VMamba adapts the Mamba state-space model, originally developed for sequence modeling in natural language processing, to the vision domain, making it particularly suitable for processing bi-temporal imagery due to its ability to capture long-range dependencies with minimal computational overhead [23]. Its promising adaptation to CD tasks is well described in [25].
As depicted in Figure 2, the VMamba encoder employs a hierarchical architecture to process input images. The process begins with a Patch Partition module that divides the input image $I^{T_{j}}$, ($j\in\{1,2\}$) into patches while embedding them into a higher-dimensional feature space. This is followed by four hierarchical stages, each consisting of $L_{i}$ Visual State-Space (VSS) blocks and, except for the first stage, a down-sampling layer. These stages generate feature maps at progressively reduced resolutions ($H/4\times W/4$, $H/8\times W/8$, $H/16\times W/16$, $H/32\times W/32$) with increasing channel depths ($C_{1},C_{2},C_{3},C_{4}$), facilitating the extraction of multi-scale features essential for both BCD and SCD.
As seen in the right panel of Figure 2, each VSS block incorporates a 2D Selective Scan (SS2D) module, which replaces the traditional self-attention mechanism found in vision transformers. The SS2D module employs a Cross-Scan strategy, scanning the 2D feature map along four directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right). This approach bridges the gap between the sequential nature of 1D state-space models and the nonsequential structure of 2D visual data, ensuring a global receptive field while maintaining linear complexity[44].
VMamba has 3 variants named VMamba-Small, VMamba-Tiny and VMamba-Base. In our implementation, we utilize the VMamba-Base variant, which outperforms other variants with $C_{1}=128,C_{2}=256,C_{3}=512,C_{4}=1024,L_{1}=2,L_{2}=2,L_{3}=15$ and $L_{4}=2$.
III-C Joint Spatio-Frequency Feature Fusion Mechanism
As illustrated in Figure 3, we propose a novel fusion mechanism $\mathcal{F}_{fusion}$ to enhance Change Detection in bi-temporal images that jointly exploits spatial cues and frequency ques to highlight subtle scene variations. At each stage $i$, the feature maps from pre-change ($X_{i}^{T_{1}}$) and post-change ($X_{i}^{T_{2}}$) images are processed as follows,
III-C1 FFT2 Branch
To capture high-frequency components, emphasizing subtle changes like edges and textures, we incorporate the FFT2 Branch, that transforms the spatial information into the frequency domain as,
$F_{i}^{T_{1}}$ $=\log(1+|\text{FFT2}(X_{i}^{T_{1}})|)$ (4)
$F_{i}^{T_{2}}$ $=\log(1+|\text{FFT2}(X_{i}^{T_{2}})|)$
where $\mathrm{FFT2}$ [^1] Implemented with torch.fft.fft2 in PyTorch, using norm=’ortho’ for orthonormal scaling. denotes the 2-dimensional (2D) Fast Fourier Transform (FFT) [50] , which operates channel-wise. For a single channel $k\in\{1,2,\dots,C_{i}\}$, $X_{i,k}^{T_{j}}\in\mathbb{R}^{H_{i}\times W_{i}}$, the 2D FFT is defined as,
$\footnotesize\mathcal{F}\!\bigl{(}X_{i,k}^{T_{j}}\bigr{)}(u,v)\;=\;\frac{1}{\sqrt{H_{i}W_{i}}}\sum_{m=0}^{H_{i}-1}\sum_{n=0}^{W_{i}-1}X_{i,k}^{T_{j}}(m,n)\;e^{-\,2\pi i\!\left(\frac{um}{H_{i}}+\frac{vn}{W_{i}}\right)}$ (5)
Here, $m$ and $n$ are spatial indices, and $u=0,1,\dots,H_{i}-1$ and $v=0,1,\dots,W_{i}-1$ are the frequency indices corresponding to the spatial dimensions $H_{i}$ and $W_{i}$, respectively. The absolute value of the FFT2 is taken, and a logarithmic scaling is applied to emphasize high-frequency components while compressing the dynamic range, making the representation more robust for detecting subtle changes.
III-C2 Difference Branch
We compute the spatial difference between the pre- and post-event feature maps to highlight potential change regions as follows,
$D_{i}=|X_{i}^{T_{1}}-X_{i}^{T_{2}}|$ (6)
III-C3 Attention-Based Spatial-Frequency Fusion
The spatial features $X_{i}^{T_{1}}$, $X_{i}^{T_{2}}$, frequency features $F_{i}^{T_{1}}$, $F_{i}^{T_{2}}$, and difference feature $D_{i}$ are concatenated along the channel axis as follows:
$X_{i}^{\mathrm{cat}}=\text{Concat}\bigl{(}X_{i}^{T_{1}},F_{i}^{T_{1}},X_{i}^{T_{2}},F_{i}^{T_{2}},D_{i}\bigr{)}.$ (7)
The concatenated feature $X_{i}^{\mathrm{cat}}$ is compressed using a $1\times 1$ convolution, yielding a reduced feature map $X_{i}^{\mathrm{reduced}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$.
To enhance informative features and suppress noise, $X_{i}^{\mathrm{reduced}}$ is refined using the Convolutional Block Attention Module (CBAM) [51], as illustrated in Figure 3. First, channel attention generates a one-dimensional gate vector, which, after sigmoid [52] activation, is multiplied element-wise with $X_{i}^{\mathrm{reduced}}$ to produce the channel-refined feature $\hat{X}_{i}^{\mathrm{reduced}}$. Subsequently, spatial attention creates a two-dimensional saliency mask, which is sigmoid activated and multiplied point-wise with $\hat{X}_{i}^{\mathrm{reduced}}$, resulting in the final fused representation $X_{i}^{\mathrm{fused}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$.
Figure 4: The Architecture of Binary-Change Decoder, that refines change predictions through a top-down pathway from Stage IV to I (left panel). At each stage, fused features $X^{\text{fused}}_{i}$ are processed by a VSS block and a CBAM-based upsampling unit (right panel). Intermediate change masks $\{CM_{i}\}_{i=1}^{4}$ are extracted to be used in CGA.
III-D Binary Change Decoder
As seen in figure 4, the Binary Change Decoder employs a top-down refinement strategy. At the coarsest stage i=4i{=}4 we first fuse $X_{4}^{T_{1}}$ and $X_{4}^{T_{2}}$ using the mechanism of Section III-C, producing $X_{4}^{\mathrm{fused}}$. A VSS Block then models long-range dependencies within this tensor. The VSS output feeds a CBAM-based upsampling unit.
CBAM-based Upsampling
As illustrated in Figure 4, the CBAM-based upsampling unit employs three parallel convolutions with kernels $1{\times}1$, $3{\times}3$, and $5{\times}5$ to gather context at complementary spatial scales, a strategy shown to enlarge the effective receptive field and sharpen object boundaries during decoder up-sampling[53, 54, 55, 56]. Each branch output is batch-normalized, concatenated, and re-weighted using CBAM’s channel and spatial attention mechanisms, which adaptively suppresses background noise. The result is combined with a convolutional branch of kernel size $1\times 1$ and interpolated to align with the H, W, and C of the next stage, before being passed to the subsequent, finer stage.
For each finer scale $i\in\{3,2,1\}$, we compute $X_{i}^{\mathrm{fused}}$ by fusing the corresponding encoder features as illustrated in 3. This fused representation is added element-wise to the upsampled output from the previous stage. The sum is then fed into a VSS block to model multi-scale context, and its output is passed through the CBAM-based upsampling module to generate features for the next stage. Additionally, we extract the VSS block output at each scale as $CM_{i}$, which serves as input to the CGA module.
III-E Semantic Map Decoders
The Semantic Map Decoders follow a similar top-down, multi-scale refinement strategy as $\mathcal{F}_{\text{BCD}}$, to produce timestamp-specific semantic maps that are conditioned on the change cues generated by $\mathcal{F}_{\text{BCD}}$ as seen in figure 5. Two weight-independent decoders, $\mathcal{F}_{\text{SCD}}^{T_{1}}$ and $\mathcal{F}_{\text{SCD}}^{T_{2}}$, share an identical architecture, allowing each to learn class priors specialised to the pre- and post-change scenes, respectively (see Eq.3).
III-E1 Change-Guided Attention (CGA)
At each stage $i$, the raw encoder feature $X_{i}^{T_{j}}$, ($j\in\{1,2\}$) is first modulated by its corresponding auxiliary change map $CM_{i}$ using the Change-Guided Attention block as follows,
$\widehat{X}_{i}^{T_{j}}=X_{i}^{T_{j}}\;\odot\;\sigma(CM_{i})$ (8)
where $\sigma$ denotes the sigmoid function and $\odot$ indicates element-wise multiplication. This mechanism guides $\mathcal{F}_{\text{SCD}}$ to focus on regions where changes are likely to occur.
At stage $i=4$, we compute $\widehat{X}_{4}^{T_{j}}$ and pass it through a VSS Block. The output of the VSS Block is then fed into a CBAM-based upsampling block, preparing it for the next stage.
At each finer stage $i\in\{3,2,1\}$, we compute $\widehat{X}_{i}^{T_{j}}$ and perform element-wise addition with the upsampled output from the previous stage. The resulting sum is then passed through a VSS Block, followed by a CBAM-based upsampling block(see section III-D), to continue the decoding process.
III-F Loss Function
Our network is trained using a composite loss that integrates widely adopted cross-entropy loss[57] and a mean Intersection over Union ($\mathrm{mIoU}$) regulariser[58] , alongside our novel Separated Kappa (SeK)[31] -based loss, which explicitly rewards semantic consistency within detected change regions.
For each output $b\in\{Y_{\text{BCD}},Y^{T_{1}},Y^{T_{2}}\}$ we minimise the pixel-wise cross-entropy loss,
$\mathcal{L}_{\mathrm{CE}}^{b}=-\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\sum_{c=0}^{C_{b}-1}\tilde{Y}^{b}_{i}(c)\,\log\!\bigl{(}P^{b}_{i}(c)\bigr{)},$ (9)
where $N_{b}$ is the number of valid (non-void) pixels, $C_{b}=2$ for the binary change map and $C_{b}$ equals the number of land-cover classes for the semantic maps, $P^{b}_{i}(c)$ is the softmax probability that pixel $i$ belongs to class $c$ in branch $b$, $\tilde{Y}^{b}_{i}(c)$ is the one-hot encoded ground-truth tensor, equal to 1 if pixel $i$ belongs to class $c$, and 0 otherwise.
To address the inherent class imbalance in change detection tasks—where unchanged pixels typically dominate the dataset—and to improve boundary delineation accuracy, we incorporate an mIoU regularizer [58]. The mIoU metric provides balanced evaluation across all classes by computing the intersection over union for each class independently, thereby preventing the model from being biased toward the majority class. This regularization term has been shown to significantly improve segmentation performance[59].
Let $Q\in\mathbb{R}^{N\times N}$ represent the confusion matrix derived from predicted and Ground Truth (GT) labels, where $q_{i,j}$ denotes the count of pixels classified as class $i$ by the model but belonging to GT class $j$, for $i,j\in\{1,\cdots,N\}$,with class 1 signifying no-change. The mIoU loss can be defined as,
$\mathcal{L}_{\text{mIoU}}=-\log\bigl{(}\mathrm{mIoU}+\varepsilon\bigr{)},$ (10)
where,
$\mathrm{IoU}_{1}=\frac{q_{11}}{\sum_{k=1}^{2}q_{k1}+\sum_{k=1}^{2}q_{1k}-q_{11}},$ (11)
$\mathrm{IoU}_{2}=\frac{q_{22}}{\sum_{i=1}^{2}\sum_{j=1}^{2}q_{ij}-q_{11}},$ (12)
$\mathrm{mIoU}=\frac{1}{2}\bigl{(}\mathrm{IoU}_{1}+\mathrm{IoU}_{2}\bigr{)},$ (13)
with $\varepsilon=10^{-6}$ ensuring numerical stability.
III-F1 Seperated Kappa
The Separated Kappa (SeK) metric, an adaptation of Cohen’s $\kappa$ tailored for semantic change detection, quantifies semantic agreement between predicted and ground-truth class labels solely within regions marked as changed, effectively excluding unchanged pixels[31]. The SeK metric is formulated as,
$\mathrm{SeK}=\exp\!\bigl{(}\mathrm{IoU}_{2}-1\bigr{)}\,\frac{\hat{\rho}-\hat{\eta}}{1-\hat{\eta}}$ (14)
where,
$\hat{\rho}=\frac{\displaystyle\sum_{i=2}^{N}q_{ii}}{\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}q_{ij}-q_{11}}$ (15)
$\hat{\eta}=\frac{\displaystyle\sum_{j=1}^{N}\bigl{(}\sum_{i}q_{ij}\bigr{)}\bigl{(}\sum_{i}q_{ji}\bigr{)}}{\bigl{(}\sum_{i,j}q_{ij}-q_{11}\bigr{)}^{2}}$ (16)
III-F2 SeK-based Loss
We compute the SeK metric at each of two timestamps, $T_{1}$ and $T_{2}$, and calculate their average as follows,
$\mathrm{SeK}_{\text{avg}}=\frac{1}{2}\left(\mathrm{SeK}_{1}+\mathrm{SeK}_{2}\right).$ (17)
While SeK can take negative values when agreement is worse than chance, we consistently observed $\mathrm{SeK}_{\text{avg}}>0$ in our experiments. Nevertheless, to handle rare degenerate cases where $\mathrm{SeK}_{\text{avg}}<0$ (indicating worse-than-chance agreement that could lead to undefined logarithmic losses and unstable gradients during optimization), we apply a clipping operation as,
$\overline{\mathrm{SeK}}=\max\left(\mathrm{SeK}_{\text{avg}},0\right).$ (18)
The SeK-based loss is then defined as,
$\mathcal{L}_{\mathrm{SeK}}=-\log\left(\overline{\mathrm{SeK}}+\varepsilon\right),$ (19)
where $\varepsilon=10^{-6}$ ensures numerical stability.
III-F3 Overall Objective
The total training objective of our network is defined as,
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}^{Y_{\text{BCD}}}+\frac{1}{2}\left(\mathcal{L}_{\text{CE}}^{Y^{T_{1}}}+\mathcal{L}_{\text{CE}}^{Y^{T_{2}}}\right)+\lambda_{1}\mathcal{L}_{\mathrm{mIoU}}+\lambda_{2}\mathcal{L}_{\mathrm{SeK}},$ (20)
where $\mathcal{L}_{\text{CE}}^{Y_{\text{BCD}}}$, $\mathcal{L}_{\text{CE}}^{Y^{T_{1}}}$, and $\mathcal{L}_{\text{CE}}^{Y^{T_{2}}}$ are the cross-entropy losses for the binary change map and semantic maps at times $T_{1}$ and $T_{2}$, respectively. $\mathcal{L}_{\mathrm{SeK}}$ enforces semantic consistency within detected change regions. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the relative importance of the mIoU and SeK losses. Through experimentation, we selected $\lambda_{1}=0.15$ and $\lambda_{2}=0.3$.
Figure 5: Semantic Map Decoder Architecture. The decoder refines change predictions via a top-down pathway from Stage IV to Stage I. At each stage, encoder feature $X^{T_{j}}_{i}$ is modulated by the ChangeMap $CM_{i}$ using a Change-Guided Attention module, processed by a VSS block, and upsampled via a CBAM-based upsampling unit to produce the semantic map $Y^{T_{j}}$ for timestamp $T_{j}$.
IV Experiments
IV-A Datasets
In this study, we utilize two primary datasets for SCD. Below, we provide detailed descriptions of these datasets, emphasizing their distinct characteristics and application scenarios.
SECOND Dataset [31]: The SECOND dataset is a semantic change detection dataset consisting of 4,662 pairs of aerial images captured at a spatial resolution of 0.53 meters/pixel. Each image is of size 512 $\times$ 512 pixels. The dataset covers urban regions including Hangzhou, Chengdu, and Shanghai, with annotations across six key land cover classes: non-vegetated surfaces, trees, low vegetation, water bodies, buildings, and playgrounds. This diversity allows for comprehensive evaluation of semantic change detection methods across various urban dynamics. We adopt the standard split of 2,968 training pairs and 1,694 test pairs for our experiments.
Landsat-SCD Dataset [60] : The Landsat-SCD dataset comprises 2,425 original image pairs with a spatial resolution of 30 meters/pixel, captured over Tumushuke, Xinjiang, China, from 1990 to 2020. These images, each sized 416 $\times$ 416 pixels, are annotated for changes across four land cover classes—farmland, desert, buildings, and water bodies—associated with 10 specific semantic change types. The dataset is divided into training, validation, and testing sets with a ratio of 3:1:1, corresponding to 1,455, 485, and 485 samples, respectively.
IV-B Implementation Details
Our architecture is implemented in PyTorch and trained using the AdamW[61] optimizer with a learning rate of $1\times 10^{-4}$, weight decay of $5\times 10^{-3}$, and batch size of 4. Training is conducted for 30,000 iterations on the SECOND dataset and 50,000 iterations on the Landsat-SCD dataset. During training, we apply comprehensive data augmentations including geometric transformations such as random rotations, horizontal and vertical flips, as well as photometric augmentations involving random adjustments to saturation, contrast, and brightness to enhance robustness against illumination and color variations. The complete source code, configuration files and pre-trained models will be publicly available at https://github.com/Buddhi19/Mamba-FCS.git upon publication.
IV-C Evaluation Metrics
To assess the performance of our framework for SCD, we adopt a comprehensive set of metrics tailored to each task, widely utilized in the literature[42] . We employ Overall Accuracy (OA), Mean Intersection over Union (mIoU), Separated Kappa (SeK), and the SCD-targeted F1 score ($F_{\text{scd}}$).
OA measures the proportion of correctly classified pixels relative to the total number of pixels, derived from the confusion matrix $Q=\{q_{ij}\}$, as described in Section III-F. This is formulated as,
$\text{OA}=\frac{\sum_{i=1}^{N}q_{ii}}{\sum_{i=1}^{N}\sum_{j=1}^{N}q_{ij}}.$ (21)
Due to the dominance of the no-change class in SCD tasks, OA alone may not fully capture performance in change regions. To address this mIoU and SeK are introduced [31] to evaluate the performance of CD and Semantic Exploitation (SE) respectively.
mIoU is computed as the average IoU across the no-change and change classes, with formulations for $\mathrm{IoU}_{1}$ and $\mathrm{IoU}_{2}$ provided in Equations (11) and (12), respectively. SeK evaluates semantic agreement within changed regions, offering a targeted assessment of SCD performance. Its formulation is detailed in III-F1 , with the corresponding loss term defined in Equation (14). $F_{\text{scd}}$ quantifies segmentation accuracy in change regions, based on precision and recall for semantic classes with change annotations. This can be formulated as,
$P_{\text{scd}}=\frac{\sum_{i=2}^{N}q_{ii}}{\sum_{i=2}^{N}\sum_{j=1}^{N}q_{ij}},$ (22)
$R_{\text{scd}}=\frac{\sum_{i=2}^{N}q_{ii}}{\sum_{i=1}^{N}\sum_{j=2}^{N}q_{ij}},$ (23)
$F_{\text{scd}}=\frac{2\cdot P_{\text{scd}}\cdot R_{\text{scd}}}{P_{\text{scd}}+R_{\text{scd}}}.$ (24)
V Results and Discussion
This section presents the experimental results on benchmark datasets. We first perform qualitative and quantitative ablation studies to evaluate the impact of individual components in the proposed architecture on SECOND dataset. Next, we compare our method’s performance, both qualitatively and quantitatively, against state-of-the-art (SOTA) methods in SCD for both SECOND dataset and Landsat-SCD dataset.
TABLE I: Quantitative results of the ablation study.
Method CGA SeK $D_{i}$ $F_{i}^{T_{1}},F_{i}^{T_{2}}$ OA (%) $F_{scd}$ (%) mIoU (%) SeK (%)
w/o CGA 88.08 63.61 73.93 24.07
w/o SeK 88.43 65.09 73.67 24.71
w/o Difference Branch 88.34 65.11 73.77 24.83
w/o FFT2 Branch 87.86 64.52 73.16 24.03
Mamba-FCS 88.62 65.78 74.07 25.50
V-A Ablation Studies
Our proposed architecture, Mamba-FCS, comprises three key components, (i) Joint Spatio-Frequency Fusion Mechanism, (ii) Change-Guided Attention (CGA) Module, and (iii) SeK Loss. To evaluate their individual contributions, we conducted ablation studies on the SECOND dataset by systematically removing each component. Specifically, we assessed the model’s performance with and without CGA and with and without SeK Loss, both quantitatively and qualitatively. Within the fusion mechanism, we evaluated the FFT2 branch both quantitatively and qualitatively, and additionally the Difference branch quantitatively.
Quantitative results, are reported in Table I. Qualitative analyses for the FFT2 branch, CGA, and SeK Loss are presented in Figures 6, 7, and 8, respectively. Each figure is structured with four columns. The first displays input bi-temporal images, the second shows the Ground Truth (GT), the third presents the model’s output without the ablated component, and the fourth depicts the full model’s output. Red boxes highlight regions where each component significantly enhances performance. The results in Table I and the qualitative visualizations collectively demonstrate that each component plays a critical role in improving the model’s performance for SCD.
Figure 6: The qualitative impact of removing the FFT2 branch. Columns (from left to right) display bi-temporal inputs, ground truth, results without the FFT2 branch, and the full model. Red boxes highlight regions where frequency-domain cues enhance edge precision and suppress hallucinated changes.
V-B Effect of FFT2 Branch
Removing the FFT2 Branch results in performance drops of 1.47%, 0.91%, 1.26%, and 0.76% in SeK, mIoU, $F_{scd}$, and OA, respectively, as shown in Table I. This underscores the FFT2 Branch’s critical role in enhancing model performance. In Figure 6, rows (a) and (b) demonstrate the FFT2 Branch’s ability to recover sub-pixel edges and suppress hallucinatory changes, which are prevalent without the FFT2 Branch. In rows (c) and (d), under challenging conditions such as heavy shadows and low contrast, the FFT2 Branch effectively detects transitions from low vegetation to ground. Rows (e) and (f) show FFT2 branch reduces boundary erosion and false changes in unchanged structures, alongside improved detection of changes from buildings to non-vegetated surfaces and reduced false detections of buildings on ground in row (f).
In summary, incorporating log-amplitude frequency features into the fusion block enables precise detection of fine-grained changes that spatial convolutions alone struggle to capture in this case. These qualitative improvements align with the quantitative gains in SeK, mIoU, $F_{scd}$, and OA.
Figure 7: The qualitative impact of removing the Change-Guided Attention (CGA) module. Columns (from left to right) display bi-temporal inputs, ground truth, results without the CGA module, and the full model. Red boxes highlight regions where CGA sharpens change masks, reduces class confusion at boundaries, and mutually enhances both BCD and SCD performance.
V-C Effect of Change-Guided Attention (CGA)
Table I shows that removing the CGA module leads to performance decreases of 1.43%, 0.14%, 2.17%, and 0.54% in SeK, mIoU, $F_{scd}$, and OA, respectively. Figure 7 further illustrates the pivotal role of CGA, which interconnects the three decoders in our architecture ($\mathcal{F}_{BCD}$, $\mathcal{F}^{T_{1}}_{SCD}$, and $\mathcal{F}^{T_{2}}_{SCD}$). Rows (a) to (d) demonstrate CGA’s ability to improve change map fidelity, and suppress false changes around changed areas. Rows (e) and (f) highlight CGA’s benefits for both BCD and SCD tasks, showing that without CGA, independent operation of decoders leads to fragmented change masks and mislabeled semantics at boundaries. By facilitating information exchange among decoders, CGA enhances coordinated learning, directly contributing to the quantitative gains observed in Table I.
Figure 8: The qualitative impact of omitting the Separated Kappa (SeK). Columns (from left to right) display bi-temporal inputs, ground truth, results without SeK, and the full model. Red boxes highlight regions where omitting SeK increases false semantic labels, while including SeK reduces artifacts.
V-D Effect of SeK Loss
Removing the SeK Loss during training results in performance drops of 0.79%, 0.40%, 0.69%, and 0.19% in SeK, mIoU, $F_{scd}$, and OA, respectively, as shown in Table I. The most significant drop occurs in the SeK metric, which emphasizes class-balanced semantic change identification, confirming that SeK Loss primarily enhances semantic class accuracy. Figure 8 provides qualitative evidence supporting this observation. Rows (a) and (c) illustrate how SeK Loss reduces false occurrences of low vegetation on ground. Similarly, rows (d) and (f) show its effectiveness in mitigating false occurrences of trees on low vegetation. Row (e) demonstrates a complete elimination of false occurrences of water. Notably, SeK Loss enhances differentiation between changed and unchanged regions across all rows, slightly improving change map accuracy.
In summary, SeK Loss strengthens semantic class identification by prioritizing class-balanced change detection, reducing false changes and improving accuracy, as evidenced by quantitative metrics in Table I and qualitative results in Figure 8.
V-E Comparative Experiments
To assess our model’s performance, we benchmark against state‐of‐the‐art methods for SCD. The CNN‐based competitors include HRSCD variants (S2–S4) [39], ChangeMask [38], SSCD-1 and Bi-SRNet [29], and TED [42]. Transformer‐based baselines are SMNet [62] and ScanNet [42]. We also compare our model with Mamba-based baselines such as ChangeMamba [28].
V-E1 Quantitative and Qualitative Results
We present quantitative results for the SECOND dataset and for the Landsat-SCD dataset in Table II. For clarity, SOTA competitors are categorized into three families: (i) CNN-based models, (ii) Transformer-based models, and (iii) Mamba-based models. Qualitative comparisons are provided through segmentation maps in Fig. 9 for the SECOND dataset and Fig. 10 for the Landsat-SCD dataset.
Figure 9: Qualitative comparison on the test split of SECOND dataset. Columns show (i) bi-temporal inputs, (ii) ground truth, and predictions from HRSCD-S4, Bi-SRNet, ScanNet, ChangeMamba, and our Mamba-FCS.
SECOND Dataset
Among CNN-based models, Bi-SRNet achieves the highest performance, with an OA of 87.84%, an $F_{scd}$ score of 62.61%, a mIoU of 73.41%, and a SeK of 23.22%. The lightweight TED model follows, demonstrating that its spatial-preserving design effectively captures semantic changes without extensive temporal reasoning. Within the Transformer family, ScanNet leverages long-range spatio-temporal dependencies, improving performance to 63.66% $F_{scd}$ and 23.94% SeK, the best results outside the Mamba family. By replacing quadratic self-attention with linear state-space mixing, ChangeMamba advances the frontier, achieving 88.12% OA, 64.03% $F_{scd}$, 73.68% mIoU, and 24.11% SeK.
Building on the same backbone and integrating the aforementioned techniques, Mamba-FCS achieves the best performance among the evaluated methods, surpassing previous results across all four metrics, achieving 88.62% OA, 65.78% $F_{scd}$, 74.07% mIoU, and 25.50% SeK on the SECOND dataset.
These results are corroborated by the qualitative analysis in Fig. 9. In rows (a), (b), (d), (h), and (i), HRSCD-str4 and Bi-SRNet consistently fail to identify the correct semantic class. While ScanNet and ChangeMamba accurately detect semantic labels in these cases, Mamba-FCS outperforms them by significantly reducing false change detections and enhancing segmentation quality, closely aligning with the GT.
TABLE II: Semantic change-detection results on the SECOND and Landsat-SCD datasets (higher is better). For each dataset, best, second-best and third-best scores are highlighted in red, blue and green, respectively. Standard train-test splits were used for both datasets.
SECOND Landsat-SCD
Method OA F1 mIoU SeK OA $F_{\text{scd}}$ mIoU SeK
CNN-based models
HRSCD-S2[39] 85.49 49.22 64.43 10.69 86.06 36.52 74.92 2.89
HRSCD-S3[39] 84.62 51.62 66.33 11.97 91.47 75.86 79.79 35.57
HRSCD-S4[39] 86.62 58.21 71.15 18.80 92.17 77.37 81.07 38.09
ChangeMask[38] 86.93 59.74 71.46 19.50 92.93 79.74 81.46 40.50
SSCD-1[29] 87.19 61.22 72.60 21.86 93.20 80.53 81.89 41.77
Bi-SRNet[29] 87.84 62.61 73.41 23.22 93.80 82.01 82.94 44.27
TED[42] 87.39 60.34 72.79 22.17 94.39 83.63 84.79 48.33
Transformer-based models
SMNet[62] 86.68 60.34 71.95 20.29 94.53 84.12 85.65 51.14
ScanNet[42] 87.86 63.66 73.42 23.94 96.04 85.62 86.37 52.63
Mamba-based models
ChangeMamba[28] 88.12 64.03 73.68 24.11 96.08 86.61 86.91 53.66
Mamba-FCS 88.62 65.78 74.07 25.50 96.25 89.27 88.81 60.26
Landsat-SCD Dataset
Among CNN-based baselines, TED leads with 94.39% OA, 83.63% $F_{\text{scd}}$, 84.79% mIoU, and 48.33% SeK. Its lightweight encoder-decoder architecture demonstrates that preserving spatial details effectively captures temporal semantics in medium-resolution Landsat imagery. In the Transformer family, ScanNet, achieves 96.04% OA, 85.62% $F_{\text{scd}}$, 86.37% mIoU, and 52.63% SeK, the best results outside the Mamba family. ChangeMamba further improves performance to 96.08% OA, 86.61% $F_{\text{scd}}$, 86.91% mIoU, and 53.66% SeK.Mamba-FCS achieves the best results among the methods evaluated in our experiments, achieving 96.25% OA, 89.27% $F_{\text{scd}}$, 88.81% mIoU, and 60.26% SeK on the Landsat-SCD dataset.
Figure 10: Qualitative comparison on the test split of Landsat-SCD dataset. Columns display (i) bi-temporal inputs, (ii) ground truth, and predictions from HRSCD-S4, Bi-SRNet, SCanNet, ChangeMamba, and Mamba-FCS. Red boxes highlight regions of improved performance for reader convenience.
These findings are supported by the qualitative analysis in Fig. 10. In rows (a) and (b), OurModel captures water regions more accurately than competing models. Similarly, in rows (c) to (j), OurModel demonstrates superior class detection, minimizes false change alarms, and produces segmentation results that closely match the GT. For readers convenience we highlight the areas of improvements in Red.
Figure 11: Dataset-level from-to confusion matrices for test split of SECOND (first row) and Landsat-SCD (second row). Cell colors indicate the percentage of changed pixels. Columns (from left to right) display ground truth, SCanNet, ChangeMamba, and Mamba-FCS. Mamba-FCS accurately captures ground-truth distributions for dominant transitions while minimizing noise in rare class predictions.
V-E2 Change Analysis
To evaluate the efficacy of each model in accurately capturing both dominant and rare class transitions, we constructed dataset-level ”from $\rightarrow$ to” confusion matrices by aggregating all changed pixels across the official test splits of the SECOND and LandSat-SCD benchmarks, as presented in Figure 11.
In the SECOND dataset, GT identifies major transitions as ground $\rightarrow$ building (32.01%), low-vegetation $\rightarrow$ ground (14.60%), and low-vegetation $\rightarrow$ building (11.59%). Our proposed Mamba-FCS closely matches these key transitions, achieving corresponding values of 34.50%, 14.06%, and 14.65%, respectively. The deviations from GT are consistently within 2.5%, except for low-vegetation $\rightarrow$ building, which differs by 3.06%. Conversely, SCanNet significantly overestimates the ground $\rightarrow$ building transition, predicting it at 38.65%, while ChangeMamba marginally shifts the low-vegetation $\rightarrow$ building transition to 13.14%. For rare transitions involving classes such as tree, water, and playground (each less than 3% occurrence in GT), Mamba-FCS maintains noise levels below 0.9%. Notably, Mamba-FCS predicts water $\rightarrow$ building at 0.29% and playground $\rightarrow$ ground at 0.26%, significantly outperforming SCanNet (8.6% overall noise) and ChangeMamba (8.1% overall noise), with Mamba-FCS maintaining a total noise of only 4.2% across the ten least frequent transitions.
In the LandSat-SCD dataset, GT predominantly includes desert $\rightarrow$ farmland (62.40%), desert $\rightarrow$ water (11.95%), and farmland $\rightarrow$ desert (9.13%) transitions. Our Mamba-FCS accurately aligns with these major transitions, reporting values of 62.98%, 11.96%, and 8.34%, respectively, with minimal deviations (all within 0.79%). In contrast, SCanNet overpredicts the desert $\rightarrow$ farmland transition at 68.10%, and ChangeMamba slightly overrepresents the desert $\rightarrow$ building transition (4.40% versus GT’s 3.84%). For infrequent transitions such as building $\rightarrow$ water and water $\rightarrow$ building (each less than 1.5% occurrence in GT), Mamba-FCS effectively controls noise levels at 0.20%, outperforming ChangeMamba (0.44%) and SCanNet (0.50%). Across all rare transitions, the cumulative noise introduced by Mamba-FCS is 2.1%, notably lower than the 3.7% from ChangeMamba and 3.9% from SCanNet.
These results demonstrate that the proposed Mamba-FCS robustly captures major land cover transitions while effectively mitigating noise from minor, infrequent class changes, underscoring its suitability and reliability for environmental monitoring and urban growth analysis tasks.
VI Conclusions
In this paper, we presented Mamba-FCS, a SCD framework specifically designed to effectively address the challenges of capturing long-range contextual dependencies and detecting subtle semantic transitions in high-resolution remote sensing imagery. Our approach combines the efficiency and global contextual modeling capabilities of VMamba, a linear-complexity state-space backbone, with an novel Joint Spatio-Frequency Feature Fusion strategy that incorporates log-amplitude frequency domain features to mitigate illumination-related artifacts and enhance fine-grained boundary detection. Additionally, we integrated a lightweight Change-Guided Attention (CGA) mechanism, which aligns semantic prediction heads with binary change cues, thereby improving semantic accuracy. Furthermore, we proposed the Separated Kappa(SeK) loss function, repurposing the established SCD evaluation metric into an effective training objective, particularly benefiting minority class transitions.
Extensive experiments on the SECOND and Landsat-SCD benchmarks demonstrate that Mamba-FCS outperforms recent methods across multiple metrics, including overall accuracy, F-score, mean Intersection over Union (mIoU), and SeK. Notably, our framework excels in detecting rare and subtle semantic transitions, achieving reduced false positives and improved class-wise accuracy. These results underscore the efficacy of Mamba-FCS in addressing complex SCD challenges in remote sensing.